OpenStudy (anonymous):

Solve, the following system of equations by linear combination: 2d + e = 8 d - e = 4

OpenStudy (ash2326):

add the two equations 2d+e+(d-e)=8+(4) 3d=12 d=4 use this in first equation 2d+e=8 d=4 2*4+e=8 e=0 so d=4 and e=0